Sunshine Trading Company (STC) has an expected EBIT of $30,000 in perpetuity, and a tax rate of 40%. Interest rate on STC’s debt is 8%, and its return on equity is 13.12%. Debt-equity ratio of STC is 0.6, or 3/5. Blackstone Inc (BI) is an all-equity firm, which is identical to STC, except for the capital structure.

(1) What is the value of STC?                             

(2) What is the value of BI?                                

(3) What is the rate of return on BI shares?
Answer #1

Sol:

1.

STC

Given , EBIT = $30000

Less Interest @ 8% = 2400

EBT 27600

Less Tax @ 40% = 11040

      

EAT 16560

Return on equity(ROE) = 13.12%

ROE = Net Income

Shareholder's Equity

or, 13.12% = 16560

Shareholder's Equity

Shareholder's Equity ( Equity Capital) = 16560

13.12%

= $ 126,220

Debt /Equity =0.6

Debt /126220 =0.6

Debt = 0.6 *126220

= $ 75732

Total capital = equity capital + debt capital

= $ 126220 +$ 75732

= $ 201952

Hence , the value of STC = $ 201952

2. Value of BI (Blackstone INC)

Since the company BI is identical to STC , it means BI is similar to STC

Here, All capital are equity only

Total capital = Equity capital + Debt capital

= $ 201952 +0

= $ 201952

Therefore, the value of BI = $ 201952

3. Rate of return on BI shares

(RR) = Kd *( 1-tax) * debt capital   + Ke  * Equity capital

Total capital Total capital

= 0 + 13.12% * 201952

201952

=13.12%

Hence , the rate of return equals to equity rate , if there is  no debt.